# Style-Preserving Lip Sync via Audio-Aware Style Reference

### v1

@@ II-A Audio-Driven Lip Sync
With the increasing application in multimedia, audio-driven lip sync has attracted widespread attention in recent years. According to the training paradigm, methods for audio-driven lip sync can be categorized into subject-specific and subject-generic. Subject-specific methods[18, 21, 22, 23, 24, 25] need to collect data of the target subject for subject-specific training. Although they can achieve lip sync conforming to the individual’s speaking style, they can not be generalized to unseen subjects without further training, thus having limited application in the real world. Related to ours, AE-NeRF [18] observes that similar pronunciations have similar lip shapes. However, instead of aggregating the reference lip shapes to model the speaking style, it aggregates the reference image features to build a robust prior for Neural radiance field (NeRF) [26] and can not generalize across different subjects without further training.
In contrast, subject-generic approaches for lip sync, such as  [5, 6, 7, 8, 9, 10, 16, 1, 2], have garnered more attention for their ability to infer lip sync for unseen subjects without further training. However, most subject-generic methods [7, 8, 9, 10, 2, 16, 1] neglect to model the speaking styles of individuals in the absence of reference information regarding individual speaking style. For example, Wav2Lip [7] pre-trains a lip synchronization discriminator on a large-scale audio-visual dataset to guide the training of a lip sync generator. Yet, it can only achieve lip sync matching the general speaking styles learned from the audio-visual dataset. Similarly, IP-LAP [8] proposes a two-stage talking face generation method based on landmark representation, using a transformer-based model to predict lip landmarks from input audio. However, due to the lack of reference information on individual speaking style, it can not generate lip motions that preserve the speaking style well.
In this paper, we propose a method that further improves subject-generic methods, achieving more accurate lip sync that better preserves the individual’s speaking style.
@@ II-B Style-Guided Lip Sync
Recent advancements in subject-generic methods, e.g. [6, 5, 11, 12, 13, 17], have made efforts to model the speaking styles of individuals through the integration of style reference videos. Among these, StyleAvatar [13] constructs a video-level style code from a style reference video based on the standard deviation of facial movements. Then, it proposes a latent-style-fusion model to synthesize stylized talking faces for arbitrary audio, imitating the talking style from the style code. Recently, StyleTalk [6] introduces a framework for generating stylized one-shot talking heads. This framework encodes the expression parameters from all frames within a style reference video into style vectors. These vectors are then aggregated into a single video-level style code through a self-attention pooling layer. Specifically, this layer calculates the attention weight for each frame by utilizing a feed-forward network, subsequently aggregating the style vectors weighted by their respective attention weights into a video-level style code. The derived style code is then used to guide the lip sync for arbitrary input audio.
Following StyleTalk [6], some later works [5, 11, 12, 17] adopt a similar style aggregation method to obtain a compact video-level style code. Despite their efforts, a common limitation persists across these methods: the inability to well preserve the speaking style of the individual during the task of lip sync. The main reason for this challenge is that, to predict the lip shape for different input audio, the network needs to reference different lip shapes from the style reference video. In response to this challenge, our study proposes a novel audio-aware style reference scheme to preserve the speaking style better for improved lip sync.
II-C Diffusion Models
Diffusion models [27] have emerged as a significant advancement in generative tasks. Compared to generative adversarial networks (GANs) [28], diffusion models exhibit enhanced training stability and alleviate the mode collapse issue, achieving superior generation quality. Particularly, latent diffusion models (LDM) [20] apply the diffusion model in the latent space of a pre-trained autoencoder to reduce the computational cost while retaining the generation quality. A further innovation is proposed in Denoising Diffusion Implicit Models (DDIM) [29], which introduce a class of non-Markovian diffusion processes to accelerate the inference process of diffusion models.
Among the applications of diffusion models, DiffTalk [10] designs a conditional latent diffusion model for lip sync tasks but overlooks modeling the speaking styles of individuals. It incorporates input audio through cross-attention layers to facilitate lip sync. Moreover, it concatenates the reference image’s latent with the noisy latent to provide subject appearance information for the latent diffusion model. In this work, we propose an advanced conditional latent diffusion model to generate lip-synced and realistic talking face videos, conditioned on the lip motion and a reference facial image.
@@ IV-A2 Implementation Details
We train all modules using the Adam optimizer [41]. The lip motion prediction network is trained with a batch size of 256 and a learning rate of 2e-4 on 4 NVIDIA 3090 GPUs. In our latent diffusion framework, the autoencoder is initialized with the pre-trained weight from latent-diffusion [42] and is frozen during training. The latent diffusion model is trained with a batch size of 12 and a learning rate of 4e-5 on 4 NVIDIA 3090 GPUs.
@@ IV-A4 Comparison Methods
In this work, we compare the proposed method with several representative state-of-the-art (SOTA) subject-generic approaches in audio-driven talking face video generation, including StyleTalk [6], DiffTalk [10], PD-FGC [16], IP-LAP [8], PC-AVS [9], and Wav2Lip [7]. StyleTalk [6] is a representative subject-generic method capable of modeling speaking style. However, the code released for StyleTalk is incomplete and unable to generate results, lacking the phoneme label extraction and training code. Consequently, we made some minor modifications to its code for comparison purposes. The modified version utilizes DeepSpeech [30] audio features instead of phoneme labels and is trained using the same objectives as ours. Additionally, StyleTalk’s image rendering network can only accept a reference facial image to provide subject appearance information, whereas other comparison methods input more conditions, such as the upper-half face at the current time step. For a fair comparison, we replaced StyleTalk’s image renderer with our diffusion rendering model to generate results. We only compare with StyleTalk in terms of lip sync quality. The methods including DiffTalk[10], IP-LAP [8], Wav2Lip[7], and Ours all synthesize talking face videos through inpainting the lower half of the face influenced by the input audio and reference facial image. Therefore, for quantitative comparison, the lower half of the face in the input video is masked and then these methods reconstruct the masked area. The input video before masking serves as the ground truth for metric computation. We train DiffTalk [10] using the released official code until convergence, but it generates temporally unstable lip motion. It relies on additional frame-interpolation techniques to smooth the results, affecting the comparison fairness. Therefore, the frame-interpolation post-processing was not utilized for a fair comparison. PC-AVS [9] utilizes a pose source video, an audio input, and a reference facial image to synthesize a talking face video. For a fair comparison, we replace its pose source video with the ground-truth video. PD-FGC  [16] inputs a pose source video, an expression source video, an eye blink source video, audio, and a reference facial image to synthesize a talking face video. For a fair comparison, we replace its pose source, expression source, and eye blink source videos with the ground-truth video.
@@ IV-B Quantitative Evaluation
TABLE I: Quantitative comparisons between the proposed StyleSync and other SOTA methods. * indicates that due to the incomplete released code of StyleTalk, we make several modifications during re-implementation. † indicates that the SyncScore only assesses the lip sync quality in terms of the general speaking style and is not suitable as a primary metric. Please see Section IV-A3 for more details. $\uparrow$ indicates higher is better, and $\downarrow$ indicates lower is better.
@@ IV-B1 Lip Sync Quality
As detailed by the LipLMD metric, the proposed method outperforms all of the comparison baselines on both VoxCeleb and HDTF datasets. This demonstrates its generation of more accurate lip sync that better preserves the speaking styles of the individuals. Specifically, Wav2Lip, PC-AVS, DiffTalk, IP-LAP, and PD-FGC lag significantly behind our proposed method in LipLMD due to their incapability to model the speaking styles of individuals. StyleTalk incorporates the use of style reference videos to model speaking styles. However, it still lags behind our StyleSync because of its inaccuracies in style aggregation. Specifically, StyleSync’s LipLMD obtains lower scores than StyleTalk by 8.4% on VoxCeleb and 13.6% on HDTF.
In SyncScore, Wav2Lip and PC-AVS achieve higher scores than the proposed method. This is because Wav2Lip utilizes SyncNet as the lip sync expert to guide the training of its generator, and PC-AVS adopts an audio-visual contrastive learning mechanism similar to SyncNet’s. However, this only indicates that their generated lip shapes are closer to the lip shapes matching the general speaking style rather than the individuals’ speaking styles, since the SyncNet for computing SyncScore overlooks modeling the speaking styles of individuals.
@@ IV-C Qualitative Evaluation
Figure 3: Qualitative comparisons between the proposed StyleSync and other SOTA algorithms, with subjects from the VoxCeleb (Left) and HDTF (Right) datasets.
For a comprehensive qualitative evaluation, we visualize some representative comparison results conducted on HDTF [14] and VoxCeleb [40] datasets in Figure 3 and give some analyses in the following subsections. To facilitate comparisons, we present the cropped faces from the generated results of all methods including the ground-truth face. Moreover, we conduct a user study from volunteers and provide a supplementary video for further evaluation.
@@ IV-C4 Supplementary Video
We have submitted a supplementary video in the review system. The supplementary video demonstrates the comparison results of all methods in several subjects from VoxCeleb [40] and HDTF [14] datasets. Please download and watch it. If there are any issues downloading the video from the review system, the same file can also be downloaded via this single-blind backup link.
@@ IV-D Style Aggregation Analysis
Figure 4: Visualization of the attention weight in cross-attention layer for style aggregation analysis. The red phonetic symbols represent the content of the audio. If a reference lip shape corresponds to the reference audio that is closer to the input audio, it has a higher attention weight, contributing more significantly to the style aggregation.
In our method, we posit that within a style reference video, if a lip shape corresponds to the reference audio that is closer to the input audio, it should contribute more significantly to the style aggregation. To verify this assumption, we visualize the attention weights of the cross-attention layers in the transformer decoder during the lip motion prediction stage. There are several cross-attention layers in the transformer decoder. We average the attention weights of all cross-attention layers to obtain a single attention weight, as shown in Figure 4. To facilitate visualization, we set the length $n$ of the style reference video to 32 frames and denote the frame numbers as $0,1,2,\cdots,31$.
As illustrated in Figure 4, the input audio corresponds to the syllable “æ”. The reference audio of the $1$-th frame in the style reference video corresponds to the syllable “e”, which is closer to the input audio’s syllable than the syllables in other frames. Therefore, the attention weight for the $1$-th frame is higher than other frames’, indicating that the reference lip shape in the $1$-th frame contributes more to the style aggregation.
@@ IV-E Ablation Study
Figure 5: Qualitative results of the ablation study. Different variants exhibit different lip sync accuracy while retaining the same visual quality. The variation of lip sync quality is more significant in the images of the third column.
@@ IV-E1 Effect of Audio-Aware Style Aggregation
The main contribution of the proposed method is the introduction of the novel audio-aware style reference scheme, which aggregates speaking style information by exploring the relationship between the input audio and the reference audios. Actually, the effectiveness of this has been verified by the fact that the LipLMD score of our method is significantly lower than that of StyleTalk as shown in Table I.
Furthermore, to verify the role of the reference audio, we devise a model variant where the reference audios from the style reference video are not utilized for style aggregation like previous methods [6, 13]. Specifically, the Transformer encoder for encoding the reference audio is removed from the proposed framework, and the reference lip features encoded from the reference lip shapes act as both keys and values for the cross-attention layers in transformer decoder. We report the result of this variant in the row "Ours w/o Reference Audio", and its LipLMD score increases by 6.5% compared to our full model’s. Besides, as shown in Figure 5, without the reference audio, the lip sync quality gets worse. This verifies the necessity to explore the relationship between reference and input audio for accurate and style-preserving lip sync.
@@ V CONCLUSION
In this paper, we have introduced a novel method called StyleSync to tackle the problem of audio-driven lip sync. It can effectively respond to the challenges of traditional methods that can only achieve sub-optimal lip sync conforming to the general speaking styles or struggle to preserve the speaking styles of individuals. Our proposed audio-aware style reference scheme, leveraging the relationships between input audio and reference audio from style reference video, effectively achieves style-preserving lip sync. In detail, the components developed by StyleSync include a novel Transformer-based lip motion prediction model that integrates speaking style information from a style reference video to accurately predict lip motions, alongside a novel conditional latent diffusion model that translates these predicted motions into realistic, lip-synced videos. These significantly contribute to producing realistic and high-fidelity talking face videos. Experiments as well as comprehensive ablation analysis have revealed StyleSync’s superiority in handling the tasks of audio-driven lip sync.

### v2

@@ II-A Audio-Driven Lip Sync
With the increasing application in multimedia, audio-driven lip sync has attracted widespread attention in recent years. According to the training paradigm, methods for audio-driven lip sync can be categorized into subject-specific and subject-generic. Subject-specific methods[18, 21, 22, 23, 24] need to collect data of the target subject for subject-specific training. Although they can achieve lip sync conforming to the individual’s speaking style, they can not be generalized to unseen subjects without further training, thus having limited application in the real world. In contrast, subject-generic approaches for lip sync, such as  [5, 6, 7, 8, 9, 10, 16, 25, 26], have garnered more attention for their ability to infer lip sync for unseen subjects without further training. However, most subject-generic methods [7, 8, 9, 10, 2, 16, 1] neglect to model the speaking styles of individuals in the absence of reference information regarding individual speaking style. For example, Wav2Lip [7] pre-trains a lip synchronization discriminator on a large-scale audio-visual dataset to guide the training of a lip sync generator. Yet, it can only achieve lip sync matching the general speaking styles learned from the audio-visual dataset. Similarly, IP-LAP [8] proposes a two-stage talking face generation method based on landmark representation, using a transformer-based model to predict lip landmarks from input audio. However, due to the lack of reference information on individual speaking style, it can not generate lip motions that preserve the speaking style well. In this paper, we propose a method that further improves subject-generic methods, achieving more accurate lip sync that better preserves the individual’s speaking style.
@@ II-B Style-Guided Lip Sync
Recent advancements in subject-generic methods, e.g. [6, 5, 11, 12, 13, 17], have made efforts to model the speaking styles of individuals through the integration of style reference videos. Among these, StyleAvatar [13] constructs a video-level style code from a style reference video based on the standard deviation of facial movements. Then, it proposes a latent-style-fusion model to synthesize stylized talking faces for arbitrary audio, imitating the talking style from the style code. Recently, StyleTalk [6] introduces a framework for generating stylized one-shot talking heads. This framework encodes the expression parameters from all frames within a style reference video into style vectors. These vectors are then aggregated into a single video-level style code through a self-attention pooling layer. Specifically, this layer calculates the attention weight for each frame by utilizing a feed-forward network, subsequently aggregating the style vectors weighted by their respective attention weights into a video-level style code. The derived style code is then used to guide the lip sync for arbitrary input audio. Following StyleTalk [6], some later works [5, 11, 12, 17] adopt a similar style aggregation method to obtain a compact video-level style code. Despite their efforts, a common limitation persists across these methods: the inability to well preserve the speaking style of the individual during the task of lip sync. The main reason for this challenge is that, to predict the lip shape for different input audio, the network needs to reference different lip shapes from the style reference video. In response to this challenge, our study proposes a novel audio-aware style reference scheme to preserve the speaking style better for improved lip sync.
@@ IV-A4 Comparison Methods
In this work, we compare the proposed method with several representative state-of-the-art (SOTA) subject-generic approaches in audio-driven talking face video generation, including StyleTalk [6], DiffTalk [10], PD-FGC [16], IP-LAP [8], PC-AVS [9], and Wav2Lip [7]. StyleTalk [6] is a representative subject-generic method capable of modeling speaking style. However, the code released for StyleTalk is incomplete and unable to generate results, lacking the phoneme label extraction and training code. Consequently, we made some minor modifications to its code for comparison purposes. The modified version utilizes DeepSpeech [27] audio features instead of phoneme labels and is trained using the same objectives as ours. Additionally, StyleTalk’s image rendering network can only accept a reference facial image to provide subject appearance information, whereas other comparison methods input more conditions, such as the upper-half face at the current time step. For a fair comparison, we replaced StyleTalk’s image renderer with our diffusion rendering model to generate results. We only compare with StyleTalk in terms of lip sync quality. The methods including DiffTalk[10], IP-LAP [8], Wav2Lip[7], and Ours all synthesize talking face videos through inpainting the lower half of the face influenced by the input audio and reference facial image. Therefore, for quantitative comparison, the lower half of the face in the input video is masked and then these methods reconstruct the masked area. The input video before masking serves as the ground truth for metric computation. We train DiffTalk [10] using the released official code until convergence, but it generates temporally unstable lip motion. It relies on additional frame-interpolation techniques to smooth the results, affecting the comparison fairness. Therefore, the frame-interpolation post-processing was not utilized for a fair comparison.
@@ IV-B Quantitative Evaluation
TABLE I: Quantitative comparisons between the proposed StyleSync and other SOTA methods. * indicates that due to the incomplete released code of StyleTalk, we make several modifications during re-implementation. † indicates that the SyncScore only assesses the lip sync quality in terms of the general speaking style and is not suitable as a primary metric. $\uparrow$ indicates higher is better, and $\downarrow$ indicates lower is better.
@@ IV-B1 Lip Sync Quality
As detailed by the LipLMD metric, the proposed method outperforms all of the comparison baselines on both VoxCeleb and HDTF datasets. This demonstrates its generation of more accurate lip sync that better preserves the speaking styles of the individuals. Specifically, Wav2Lip, PC-AVS, DiffTalk, IP-LAP, and PD-FGC lag significantly behind our proposed method in LipLMD due to their incapability to model the speaking styles of individuals. StyleTalk incorporates the use of style reference videos to model speaking styles. However, it still lags behind our StyleSync because of its inaccuracies in style aggregation. Specifically, StyleSync’s LipLMD obtains lower scores than StyleTalk by 8.4% on VoxCeleb and 13.6% on HDTF. In SyncScore, Wav2Lip and PC-AVS achieve higher scores than the proposed method. This is because Wav2Lip utilizes SyncNet as the lip sync expert to guide the training of its generator, and PC-AVS adopts an audio-visual contrastive learning mechanism similar to SyncNet’s. However, this only indicates that their generated lip shapes are closer to the lip shapes matching the general speaking style rather than the individuals’ speaking styles, since the SyncNet for computing SyncScore overlooks modeling the speaking styles of individuals.
@@ IV-C Qualitative Evaluation
Figure 3: Qualitative comparisons between the proposed StyleSync and other methods, with subjects from the VoxCeleb (Left) and HDTF (Right) datasets.
For a comprehensive qualitative evaluation, we visualize some representative comparison results in Figure 3. To facilitate comparisons, we present the cropped faces from the generated results of all methods including the ground-truth face. Moreover, we conduct a user study from volunteers and provide a supplementary video for further evaluation.
@@ IV-C4 Supplementary Video
We have submitted a supplementary video in the review system. If there are any issues downloading the video from the system, the same file can also be downloaded via this single-blind backup link.
@@ IV-D Style Aggregation Analysis
Figure 4: Visualization of the attention weight in cross-attention layer for style aggregation analysis. The red phonetic symbols represent the content of the audio. If a reference lip shape corresponds to the reference audio closer to the input audio, it receives higher attention, contributing more to the style aggregation.
In our method, we posit that within a style reference video, if a lip shape corresponds to the reference audio that is closer to the input audio, it should contribute more significantly to the style aggregation. To verify this assumption, we visualize the attention weights of the cross-attention layers in the transformer decoder during the lip motion prediction stage. There are several cross-attention layers in the transformer decoder. We average the attention weights of all cross-attention layers to obtain a single attention weight, as shown in Figure 4. To facilitate visualization, we set the length $n$ of the style reference video to 32 frames and denote the frame numbers as $0,1,2,\cdots,31$. As illustrated in Figure 4, the input audio corresponds to the syllable “æ”. The reference audio of the $1$-th frame in the style reference video corresponds to the syllable “e”, which is closer to the input audio’s syllable than the syllables in other frames. Therefore, the attention weight for the $1$-th frame is higher than other frames’, indicating that the reference lip shape in the $1$-th frame contributes more to the style aggregation.
@@ IV-E Ablation Study
Figure 5: Qualitative ablation results. Different variants exhibit different lip sync accuracy while retaining the same visual quality. The variation of lip sync quality is more significant in the images of the third column.
@@ IV-E1 Effect of Audio-Aware Style Aggregation
The main contribution of the proposed method is the introduction of the novel audio-aware style reference scheme, which aggregates speaking style information by exploring the relationship between the input audio and the reference audios. Actually, the effectiveness of this has been verified by the fact that the LipLMD score of our method is significantly lower than that of StyleTalk as shown in Table I. Furthermore, to verify the role of the reference audio, we devise a model variant where the reference audios from the style reference video are not utilized for style aggregation like previous methods [6, 13]. Specifically, the Transformer encoder for encoding the reference audio is removed from the proposed framework, and the reference lip features encoded from the reference lip shapes act as both keys and values for the cross-attention layers in transformer decoder. We report the result of this variant in the row "Ours w/o Reference Audio", and its LipLMD score increases by 6.5% compared to our full model’s. Besides, as shown in Figure 5, without the reference audio, the lip sync quality gets worse. This verifies the necessity to explore the relationship between reference and input audio for accurate and style-preserving lip sync.
@@ V CONCLUSION
In this paper, we have introduced a novel method called StyleSync to tackle the problem of audio-driven lip sync. It can effectively respond to the challenges of traditional methods that can only achieve sub-optimal lip sync conforming to the general speaking styles or struggle to preserve the speaking styles of individuals. Our proposed audio-aware style reference scheme, leveraging the relationships between input audio and reference audio from style reference video, effectively achieves style-preserving lip sync. In detail, the components developed by StyleSync include a novel Transformer-based lip motion prediction model that integrates speaking style information from a style reference video to accurately predict lip motions, alongside a novel conditional latent diffusion model that translates these predicted motions into realistic, lip-synced videos. Experiments as well as comprehensive ablation analysis have revealed StyleSync’s superiority in audio-driven lip sync.